How many grams of water vapor is produced by burning 78 g of ethine?


1
Expert's answer
2021-03-10T06:57:25-0500

1 mole of "C_2H_2" =2(12.0)+2(1.0)=26g

1 mole of "H_2O" =2(1.0)+16.0=18.0g


"=\frac{78g ×1 mole} {26g}" = 3 mol "C_2H_2"



"=\frac{3.0 mol C_2H_2×2 H_2O}{ 2 C_2H_2} =3.0 mol H_2O"


=3mol "H_2O" ×18.0g


=54g of water vapor.